OpenStudy (anonymous):

Solve the equation, 9^4x-3=27^5x

imqwerty (imqwerty):

\[9^4x-3=27^5x\] is this ur ques?

OpenStudy (michele_laino):

I think it is: \[\Large {9^{\left( {4x - 3} \right)}} = {27^{5x}}\] @imqwerty

imqwerty (imqwerty):

ah yes :)

OpenStudy (anonymous):

In the first part, the x-3 is in with the power, and in the second part the x is also in with the power. Yes, Michele_Laino that's right

imqwerty (imqwerty):

\[9^{4x-3}=27^{5x}\]\[3^{2(4x-3)}=3^{3(5x)}\]when bases are same powers are equal so 2(4x-3)=5x can u get x from here?

OpenStudy (anonymous):

So now do you just solve for x?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Does x=2?

OpenStudy (michele_laino):

hint: we can simplify the equation above like below: \[\Large \begin{gathered} 2\left( {4x - 3} \right) = 3\left( {5x} \right) \hfill \\ 8x - 6 = 15x \hfill \\ \end{gathered} \]

OpenStudy (aaronandyson):

:)

OpenStudy (anonymous):

So the answer is -6/7?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

Okay, so question with the hint that you gave me where did you get the 3(5x) from?

OpenStudy (anonymous):

Okay, so question with the hint that you gave me where did you get the 3(5x) from?

OpenStudy (michele_laino):

since I can write this: \[\huge {27^{5x}} = {\left( {{3^3}} \right)^{5x}} = {3^{\left( {3 \cdot 5x} \right)}}\]

OpenStudy (anonymous):

Ohh, okay. Got it now. Thanks so much!!

OpenStudy (michele_laino):

:)